A bag is full of marbles. You reach in and pull out a single marble. If P(red)= 0.10, P(blue)=0.05, P(green)=0.25, P(yellow)=0.30, and P(purple)= 0.30, and P(purple)=0.30, then what is the probability of selecting a red or yellow marble?
Answer #1

P(red)=0.10

P(yellow)=0.30

Thus, P(selecting a red or yellow marble)

= P( red U yellow)

= P (red) + P(yellow)   

= 0.10+0.30

= 0.40

( As we draw a marble from bag it can be of only one colour. Thus we add the probability in this case. P(red or yellow) = P(red)+P(yellow) . )